The sum of a two-digit number is 11. The difference between the number and the number with its digits reversed is 27. What is the original number?
I got 74 as the original number but I can't really figure out a formula to go with this problem.

Answer by scott8148(6628)

you have to solve for the individual digits ___ let t=tens digit and u=units digit

"The sum of a two-digit number is 11" ___ t + u = 11

"The difference between the number and the number with its digits reversed is 27" ___ 10t + u + 27 = 10u + t

9t + 27 = 9u ___ t + 3 = u

substituting ___ t + (t + 3) = 11 ___ t = 4

substituting ___ (4) + u = 11 ___ u = 7